OpenStudy (anonymous):

Find the volume of the sold described. Region R is bounded by Y=x^2/3, X=0, X= -8, and X=0 The base and cross sections are squares. 16 right point subdivisions.

OpenStudy (phi):

is that \[ y= \frac{x^2}{3} \] ? also, you have x=0 twice... I don't see a volume... just a plane figure.

OpenStudy (anonymous):

no its y= x ^ 2/3 and sorry -8,8,0 @phi